GFU Insurance Company offers mail-order car insurance to preferred-risk drivers in the New York city. The company is the cheap provider of insurance in this market but doesn’t believe its annual premium of $2,500 can be raised for competitive reasons. Rates are expected to remain stable during coming periods; hence, P = MR = $2,500. Total and marginal cost relations for the company are as follows:

TC = $100,000 + $100Q + $0.006 Q2

MC = ∂TC/∂Q = $100 + $0.012Q

The profit maximizing activity level, and the profit at this activity level would be?
1
Expert's answer
2020-12-14T05:41:07-0500

The profit maximizing activity level is:

P = MR = MC,

100 + 0.012Q = 2,500,

Q = 200,000,

"TP = TR - TC = 2,500\u00d7200,000 - (100,000 + 100\u00d7200,000 + 0.006\u00d7200,000^2) = 239,900,000."